# Disentangled Dynamic Intrusion Detection

### v2

@@ V-A Experimental Settings
Figure 6: Comparisons of multi-classification. Here $\dagger$ indicates that the results are directly copied from the previous works.
V-B Main Results
V-B1 Comparisons of binary classification
@@
Under this setting, we classify a traffic flow as an attack or a benign one. We categorize the baselines into three groups, including dynamic GCNs at the top of Table I, static GCNs at the middle of the table, and another three baselines at the bottom. It should be noted that AnomRank and DynAnom are two popular baselines for anomaly detection. We run our experiment $5$ times and report the mean and variance values. The comparison results in Table I show that our DIDS consistently performs the best among all baselines over the five benchmarks, which shows the superiority of our method for intrusion detection. Specifically, compared to the E-GraphSAGE, the previous state-of-art GCN-based approach, our method achieves a $4.80\%$ higher F1-score over the CIC-BoT-IoT dataset. Our DIDS outperforms AnomRank, the previous state-of-the-art method for anomaly detection on F1, by $15.27$ points over the EdgeIIoT dataset. We attribute the above results to the gains of our statistical disentanglement, representational disentanglement and dynamic graph diffusion method.
V-B2 Comparisons of multi-classification
@@
We compare the performance of our method to four baselines in declaring the specific attack type. These baselines include E-GraphSAGE [17], TGN [91], ML [98], and AdaBoost [99], as they are representative of different types of intrusion detection models and have been widely used in previous studies. Figure 6 shows that the proposed DIDS consistently achieves the best results among all others over five datasets. For example, ours yields higher classification accuracy of up to $25$ points compared to the baseline TGN for detecting Injection attacks. The existing graph-based methods, including E-GraphSAGE, ML, and AdaBoost, perform inconsistently in the identification of complex attacks (e.g., MITM, Uploading, and XSS). We also observe that some attacks that are not easily detected by the baseline approaches, can be identified by the proposed DIDS with high F1 scores. For example, E-GraphSAGE only achieves $18.34\%$ and $30.7\%$ F1 scores on CIC-ToN-IoT for MITM and Backdoor attacks, respectively, while our DIDS is able to obtain higher than $23\%$ F1 score for each attack. These results further show the superiority of the proposed DIDS. We also find that the average scores of our method on CIC-BoT-loT and CIC-ToN-loT are lower than the ones on the other three datasets. The underlying reason is the unbalanced attack distributions in the training set, where the dominant type may mislead the classifications. Such a finding aligns with previous work in the field of computer vision [100]. Nevertheless, the proposed DIDS is still the best under such distributions. We leave this interesting observation as our future work.
TABLE II: Comparisons of few-shot learning classification on five datasets. The results are the average scores of ten-time repetitions.
F1 NMI F1 NMI F1 NMI F1 NMI F1 NMI
MBase [22] $41.74$$\pm$$3.22$ $35.81$$\pm$$2.84$ $38.52$$\pm$$3.58$ $27.89$$\pm$$3.84$ $51.13$$\pm$ $2.31$ $55.61$$\pm$ $2.05$ $41.74$$\pm$$3.22$ $35.81$$\pm$$2.84$ $63.74$$\pm$$4.14$ $68.76$$\pm$$3.16$
MTL [62] $30.79$$\pm$$3.13$ $29.81$$\pm$$3.58$ $37.80$$\pm$$5.66$ $25.10$$\pm$$2.69$ $52.64$$\pm$ $5.69$ $55.41$$\pm$ $4.16$ $44.18$$\pm$ $4.22$ $35.85$$\pm$ $2.79$ $45.38$$\pm$ $5.88$ $47.04$$\pm$ $4.82$
  TEG [101] $29.92$$\pm$ $2.17$ $22.64$$\pm$ $2.22$ $29.62$$\pm$ $3.74$ $25.71$$\pm$ $2.82$ $26.34$$\pm$ $2.45$ $27.58$$\pm$ $2.64$ $27.80$$\pm$ $2.14$ $21.44$$\pm$ $1.74$ $25.20$$\pm$ $3.17$ $23.48$$\pm$ $3.97$
CLSA [71] $17.21$$\pm$ $3.63$ $45.01$$\pm$ $2.83$ $38.54$$\pm$ $3.49$ $51.70$$\pm$ $4.90$ $18.84$$\pm$ $3.10$ $44.25$$\pm$ $3.27$ $17.53$$\pm$ $1.95$ $45.67$$\pm$ $4.13$ $19.19$$\pm$ $2.90$ $49.76$$\pm$ $3.46$
ESPT [72] $36.76$$\pm$ $4.37$ $38.51$$\pm$ $3.29$ $39.41$$\pm$ $5.37$ $34.66$$\pm$ $3.49$ $61.28$$\pm$ $8.34$ $53.77$$\pm$ $6.91$ $46.58$$\pm$ $5.94$ $49.37$$\pm$ $6.58$ $41.37$$\pm$ $4.98$ $46.52$$\pm$ $6.71$
ICI [70] $55.67$$\pm$ $5.14$ $47.80$$\pm$ $3.77$ $71.32$$\pm$ $3.35$ $57.51$$\pm$ $4.50$ $49.74$$\pm$ $4.65$ $51.30$$\pm$ $5.27$ $39.17$$\pm$ $3.66$ $31.81$$\pm$ $2.21$ $76.53$$\pm$ $9.43$ $79.00$$\pm$ $6.20$
KSCL [73] $36.19$$\pm$ $2.66$ $38.17$$\pm$ $3.90$ $33.50$$\pm$ $4.23$ $29.40$$\pm$ $4.80$ $34.87$$\pm$ $2.14$ $30.63$$\pm$ $2.56$ $17.03$$\pm$ $1.97$ $21.90$$\pm$ $1.98$ $25.74$$\pm$ $3.01$ $28.26$$\pm$ $3.15$
BSNet [27] $58.41$$\pm$$2.36$ $57.76$$\pm$$2.96$ $42.00$$\pm$$5.19$ $31.41$$\pm$$4.71$ $43.74$$\pm$ $4.20$ $49.68$$\pm$ $5.09$ $65.73$$\pm$$2.65$ $58.96$$\pm$$2.08$ $56.39$$\pm$$5.75$ $58.56$$\pm$$3.76$
CMFSL [31] $37.46$$\pm$$5.37$ $47.51$$\pm$$4.29$ $58.62$$\pm$ $3.58$ $51.47$$\pm$ $6.28$ $39.43$$\pm$ $3.88$ $43.67$$\pm$ $5.75$ $51.30$$\pm$ $4.29$ $47.62$$\pm$ $6.47$ $61.37$$\pm$ $4.95$ $58.42$$\pm$ $6.17$
TAD [32] $49.80$$\pm$$2.13$ $42.23$$\pm$$1.35$ $62.75$$\pm$$6.40$ $56.53$$\pm$$7.70$ $50.82$$\pm$ $3.93$ $53.40$$\pm$ $3.86$ $58.97$$\pm$$3.76$ $58.72$$\pm$$2.60$ $54.88$$\pm$$3.34$ $52.15$$\pm$$1.57$
PCWPK[102] $42.92$$\pm$$3.80$ $44.17$$\pm$$3.12$ $56.08$$\pm$$2.91$ $53.01$$\pm$$5.17$ $34.74$$\pm$$2.54$ $34.29$$\pm$$3.36$ $51.97$$\pm$$1.76$ $43.62$$\pm$$2.27$ $52.39$$\pm$$8.60$ $64.56$$\pm$ $5.57$
Ours(DIDS-MFL) 97.47$\pm$1.17 94.27$\pm$2.62 96.64$\pm$1.88 91.27$\pm$3.94 92.73$\pm$2.57 88.21$\pm$3.25 97.39$\pm$1.41 94.32$\pm$2.84 93.93$\pm$3.27 90.98$\pm$3.69
V-B3 Comparisons of unknown attacks
@@
To further investigate the performance of detecting unknown attacks, we conduct experiments on the four attack types by discarding the corresponding instances in the train set and detecting them in the test set. From this perspective, the unknown attacks are the new-type attacks that do not belong to the existing types of attacks in the training set. We run our experiments $5$ times and report the mean and variance values with different random seeds. Table III reports the classification results on the CIC-ToN-IOT dataset. It shows that the statistical rule-based method Logistic Regression can only achieve as low as a $1.68$ F1 score for DDoS attacks, this confirms our analysis at the very beginning that rule-based methods can hardly detect unknown attacks. The score of graph-based E-GraphSAGE is much smaller than DIDS, e.g., $6.05\%$ for MITM, indicating the limitations of the static graph in detecting unknown attacks. We also observe that TGN performs better than E-GraphSAGE, although both of them are graph-based methods. We attribute the improvement to the dynamic module for TGN. Nevertheless, our DIDS outperforms all these methods by a large margin, with an average score of $33.65\%$ on the four attacks. The results also suggest that our method is more consistent in detecting various unknown attacks, showing the effectiveness of the two disentanglements. We further provide empirical studies on more datasets in Appendix 16 to showcase the unknown detection ability of DIDS.
TABLE III: Unknown classification on the CIC-ToN-IoT dataset with metric F1-score (%)
Method Logistic Regression MSteam E-GraphSAGE TGN Ours (DIDS)
@@
V-B4 Comparisons under few-shot settings
Task Setting: We first conduct the DIDS pretraining over the known benign and attack traffic. Then we conduct a $\mathcal{N}$-way-$\mathcal{K}$-shot learning for the few-shot attack traffic. Specifically, $\mathcal{N}$ is the category number of few-shot traffic, and $\mathcal{K}$ is the training number of each category, referred to support set. In our few-shot experiments, we set $\mathcal{K}$ as $5$. We also construct a query set for our few-shot task, with $15$ sample numbers for each category. The query set can be sampled from the category samples, or augmented from the support set. We split a $5$-fold cross-validation set from the training set, with $20$% of support and query set. Taking CIC-TON-IoT as an example, we conduct a five-way-five-shot learning and five-classification test with $100$ training samples. We repeat the above training, validate, test ten times, and report the average performance. Each time’s training and validation samples are randomly selected from our existing traffic data, thus simulating the few-shot learning setting in real-world scenarios.
Baselines: To evaluate the performance of the proposed few-shot learner, we select $11$ few-shot learning models to incorporate into DIDS as baselines, including $3$ meta-learning based models (i.e., MBase [22], MTL [62], TEG [101]), where TEG is designed for graph-structure-based few-shot learning, $4$ augmentation-based models (i.e., CLSA [71], ESPT [72], ICI [70], KSCL [73], where CLAS and ESPT are based on contrastive augmentation, ICI and KSCL are based on instance augmentation), $4$ metric learning-based models (i.e., BSNet  [27], CMFSL  [31], TAD [32], PCWPK [102]).
Performance metrics: We follow the previous work [103] using F1 scores and NMI[104] metrics for multi-classification comparison.
As shown in Table II, our MFL consistently performs better among $11$ baselines under the five benchmarks, showing the superiority of our proposed MFL module in few-shot intrusion detection. Specifically, our MFL achieves reproducible average results of $95.63$% and $91.81$% in F1-score and NMI value over five public datasets. The results gain $97.87$% - $126.99$% and $73.59$% - $139.53$% improvements in average F1-score and NMI value of $11$ baselines over five public datasets, respectively. We attribute the superior and impressive reproducible results of MFL to our multi-scale representation information using and disentanglement among few-shot traffic representations. The $RQ5$ and $RQ6$ in the discussion section further verified our attributions. Furthermore, we also discuss the practical time cost of DIDS-MFL and the superiority of our model designs in time cost in Appendix 20.
V-C Ablation Study
V-C1 DIDS
In this section, we conduct an ablation study on the CIC-ToN-IoT dataset to evaluate the effectiveness of each component. We remove our statistical disentanglement and denote it as "w/o SD". We use "w/o RD" and "w/o MLGRAND" to refer to the model that removes representational disentanglement and the multi-layer graph diffusion module, respectively. Table IV reports the comparison results. It shows that removing the multi-layer graph diffusion module leads to the most significant performance degradation, e.g., an $18.33$ points decrease in AUC, indicating that it is the key component for the accuracy of the proposed DIDS. Our second disentangled memory is also non-trivial to the overall detection accuracy, as removing this component can decrease the performance by $12.47$ points in AUC. We observe that the SD module also benefits the model performance. The above ablation study further confirms the effectiveness of the three key components.
TABLE IV: Ablation study of DIDS. F1-score (%), Precision (%), ROC-AUC (%), Recall (%), all metrics above are the average of five repeated experiments on the CIC-ToN-IoT dataset.
@@ V-C1 DIDS
DIDS(ours) 97.78$\pm$0.32 98.06$\pm$0.43 97.92$\pm$0.26 96.04$\pm$0.25
V-C2 MFL
Furthermore, we conduct an ablation study on the CIC-TON-IOT dataset to evaluate the effectiveness of our designed multi-scale few-shot learning (MFL) module. We denote our model without multi-scale latent optimization space as "w/o LOS" and "w/o DR" as our model without the disentanglement regularization term. For the model without both of the above components, we denote it as "SE", which degenerated into a naive self-expressiveness model. Table V shows that removing both LOS and DR leads to dramatic performance degradation, i.e., $9.92$% and $9.75$% drop in F1-score and NMI, respectively, indicating the effectiveness of our designed MFL. Furthermore, we observed that each component of MFL, e,g. LOS or DR is necessary for performance improvement, cause removing one of them will lead to performance degradation, i.e., $22.21$%-$23.16$% and $8.63$%-$9.24$% drop in F1-score and NMI, respectively. Without any one of these two components, the performance will be poorer than the base model SE. Results showcase the complementary relationships of the proposed two modules. The LOS provides a latent optimization space across multi-scale representations, which is the foundation of the DR term. Then the DR term generates effective disentangled representations across multiple scales based on LOS.
TABLE V: Ablation study of MFL. F1-score (%), Precision (%), ROC-AUC (%), Recall (%), all metrics above are the average of ten repeated experiments on the CIC-ToN-IoT dataset.
Variants P R F1 NMI
w/o LOS $77.80$$\pm$$3.83$ $78.80$$\pm$$1.99$ $75.82$$\pm$$2.89$ $86.13$$\pm$$2.15$
w/o DR $75.51$$\pm$$3.01$ $77.30$$\pm$$1.77$ $74.90$$\pm$$1.66$ $85.56$$\pm$$1.64$
SE $88.62$$\pm$$3.56$ $88.80$$\pm$$3.27$ $87.80$$\pm$$3.17$ $85.08$$\pm$$2.31$
MFL(ours) 97.68$\pm$1.09 97.50$\pm$1.18 97.47$\pm$1.17 94.27$\pm$2.62
V-D Discussion
DIDS:
@@ V-D Discussion
Figure 7: Statistical disentanglement of traffic features.
@@
RQ2222: How does the representational disentanglement benefit "highlighting the attack-specific features"? To answer this question, we track several Injection attack data in the CIC-ToN-IoT dataset and obtain the representation of these data in DIDS and E-GraphSAGE. Meanwhile, we calculate the above two methods’ average values of embeddings for the benign traffic. As shown in Figure 8, the representation values of E-GraphSAGE are much closer to the normal. It illustrates that as nodes aggregate, the discrepancies in features become blurred, leading to inaccurate classification. While benefiting from the representational disentanglement, each dimension of features in DIDS can effectively preserve its own properties, deviating from the averages. Especially for the attack-specific features Fwd Pkt Len Max and Fwd Pkt Std, etc., these attack-specific features are significantly highlighted in Fig. 1(a), thus improving the accuracy of detection. The result proves the effectiveness of the proposed DIDS in maintaining a disentangled representation during the aggregation process, ensuring the presence of discrepancies, thus highlighting the attack-specific features and leading to more accurate classification for attacks.
(a) DIDS
(b) E-GraphSAGE
@@
Figure 8: The comparison of node representation of the Injection attack after graph aggregation of our DIDS and E-GraphSAGE. The grey line presents benign data.
@@
RQ3333: How does the multi-layer diffusion module perform effectively for intrusion detection? We have illustrated the principle of multi-layer diffusion in Section IV-C. In this part, we take the MITM attack as an example to illustrate the effectiveness of spatial-temporal in intrusion detection.
@@
Figure 9 (a) shows a deep learning-based NIDS. When a MITM attack occurs, it is difficult to detect the intrusion since the spatial and temporal information of those packets is not considered. There are also some methods that only consider a single aspect of spatial and temporal information, such as E-GraphSAGE and MStream. In this case, for example, E-GraphSAGE mainly focuses on the spatial relationship of the set nodes and extracts features from them. However, we observe that different streams have their own timestamps from Table VI, so the lack of temporal information makes it impossible to analyze the dynamic structural changes of the edge. Similarly, taking the temporal information as the only effect factor will also get incomplete characteristics that do not contain spatial information (IP address). Moreover, some methods that take both the spatial and temporal information into account, such as Euler, take the snapshot method to capture the feature of the flow which does not achieve the synchronous update for spatial and temporal information. As shown in Figure 9 (b), intuitively, we can quickly detect that UE6 is an intrusion device of layer 1 when the flow changes from $SW2-SW3$ to $SW2-UE6$ and $UE6-SW3$ considering $SW2$, $SW3$ are layer $2$ devices. Also, we have noticed the changes in dynamic graph structure with a multi-Layer graph diffusion module to realize spatio-temporal coupling and synchronous updating. Overall, DIDS performs best among these baselines in detecting various attacks.
TABLE VI: TimeStamp and IP
@@
Figure 9: Spatial-temporal coupling in intrusion detection.
@@
RQ4444: How does the disentanglement facilitate the explainability of DIDS? For this question, we rely on Figure 8 (a) as an example to recover the possible traffic features of a password attack. Since the original features are retained after the disentanglement, we can find some feature values that deviate significantly from the normal values in the node embedding. In the password attack, we observe that the deviated features after disentanglement are "Fwd Pkt Std" and "Fwd Pkt Len Max". It aligns with our common sense for the main causes of password attacks and further benefits the explainability of DIDS.
MFL:
RQ5555: How does multi-scale transform-based MFL benefit the distinction of few-shot traffic threats? To answer this question, we visualize the learned representations of DIDS and the few-shot learning module MFL on the CIC-TON-IoT dataset via t-SNE technology. As shown in Fig. 10(a) (b), the MFL’s representations are highly separated and distinguishable for different few-shot attacks, i.e., the red square box, compared to the representation generated by DIDS, i.e., the red round box. It verifies the effectiveness of our proposed multi-scale few-shot learning framework, i.e., multiple coefficient matrices fusion and multi-scale transformation. They discover the attack-specific invariant features among few-shot traffic in latent space, thus improving the distinction of attack representations.
(a) The t-SNE of DIDS
(b) The t-SNE of MFL
Figure 10: The t-SNE visualization of representations on the CIC-TON-IoT dataset generated by DIDS and MFL.
RQ6666: Can MFL disentangle the representations of few-shot samples? To answer this question, we visualize the learned representations of DIDS and MFL on the CIC-TON-IoT dataset via correlation heatmaps. As shown in Fig. 11(a) (b), the visualization results significantly reveal that MFL can generate highly disentangled representations via our designed regularization term. Specifically, MFL generates a block diagonal heatmap with high correlations and non-diagonal areas with very low correlations. It verifies that MFL disentangles the few-shot traffic representations and highlights the attack-specific ones, making them more distinguishable.
(a) The correlation map of DIDS
(b) The correlation map of MFL
Figure 11: The correlation map and the t-SNE visualization of representations generated by DIDS and DIDS-MFL with MFL.
RQ7777: Can MFL be served for other intrusion detection methods to strengthen their few-shot learning ability? To answer this question, we use MFL as a plug-and-play module to incorporate into other intrusion detection methods, i.e., E-GraphSAGE, EULER, and TGN. We follow our few-shot setting to conduct experiments on these methods with and without MFL. The methods with MFL significantly outperform the ones without MFL by $17.45$%-$544.77$%, as shown in Table VII. It demonstrates the effectiveness of our proposed MFL as a plug-and-play module to strengthen the few-shot learning ability of other NIDS. Additionally, the underlying cause of the slight improvement of E-GraphSAGE and EULER with MFL is interesting for future few-shot NIDS designs. The performance differences between E-GraphSAGE and TGN lie in the generated representation qualities of respective approaches, with more details available in Appendix 18. We also provide the visualizations of different baselines to showcase the disentangle representation ability of MFL in Appendix 19.
TABLE VII: The few-shot learning F1-score (%) comparison between different methods with and without MFL on the CIC-TON-IoT dataset.
Methods E-GraphSAGE EULER TGN
Original $20.14$$\pm$$0.15$ $52.38$$\pm$$0.29$ $14.25$$\pm$$0.33$
MFL 26.70$\pm$0.21 61.52$\pm$0.37 91.88$\pm$0.14
RQ8888: Can MFL improve the performance of DIDS? So far, we have verified the effectiveness of MFL in few-shot traffic intrusion detection, including serving as a plug-and-play module for other methods. To further study the applicability of MFL in large-scale traffic, we conduct DIDS training with MFL and report the multi-classification results, as shown in Table VIII. DIDS with MFL achieves a $2.31$% to $15.82$% F1-score improvements among three datasets. The results reveal the effectiveness of MFL when serving as a multi-classification module for normal-size traffic training. Our designed multi-scale transform-based framework also sheds light on the future intrusion detection model design.
TABLE VIII: The multi-classification F1-score (%) comparison between different datasets on DIDS with MFL module.
Datasets CIC-TON-IoT CIC-BoT-Iot NF-UNSW-NB15-v2
DIDS $83.56$$\pm$$1.18$ $88.34$$\pm$$1.75$ $93.55$$\pm$$0.97$
DIDS-MFL 96.78$\pm$1.72 91.22$\pm$3.45 95.71$\pm$5.89
RQ9999: Network intrusion detection and Large language model (LLM). Recently, there has been a significant surge in the development and application of Large Language Models (LLMs) [105]cross various domains, including Natural Language Processing (NLP)[106], Computer Vision (CV)[107], and multimodal tasks. However, the existing LLMs, e.g. GPT-$3.5$, GPT-$4$ struggle to detect traffic threats accurately and efficiently. As shown in Table IX, the F1-score of the existing LLM is significantly lower than our SOTA method DIDS. It may lie in the hallucinations[108] of existing neural language-based LLM, which struggle to comprehend the intrinsic attack features in traffic data. Second, the LLM necessitates more time to process the input traffic data and may fail as the size of the input traffic increases. It may be attributed to the numerical values of the traffic data involved. To sum up, the future directions to empower NIDS via LLM can be: 1. Aligning encrypted traffic data with the input requirements of LLMs, thus speeding the intrusion detection via LLM; 2. Determining the necessary volume of traffic data and corresponding traffic mining technology to unlock the powerful inference capabilities of LLM.
TABLE IX: The F1-score (%) and time cost (s) comparisons between the LLM and DIDS-MFL on 100 traffic samples from the CIC-TON-IoT dataset.
F1-Score Time Cost
LLM $66.67$% $13.88$s
DIDS-MFL 99.52% 6.41s
VI CONCLUSION

### v3

@@ 5.1. Experimental Settings
Figure 4. Comparisons of multi-classification. Here $\dagger$ indicates that the results are directly copied from the previous works.
5.2. Main Results
5.2.1. Comparisons of binary classification.
@@ 5.2. Main Results
Under this setting, we classify a traffic flow as an attack or a benign one. We categorize the baselines into three groups, including dynamic GCNs at the top of Table 1, static GCNs at the middle of the table, and another three baselines at the bottom. It should be noted that AnomRank and DynAnom are two popular baselines for anomaly detection. We run our experiment $5$ times and report the mean and variance values. The comparison results in Table 1 show that our 3D-IDS consistently performs the best among all baselines over the five benchmarks, which shows the superiority of our method for intrusion detection. Specifically, compared to the E-GraphSAGE, the previous state-of-art GCN-based approach, our method achieves a $4.80\%$ higher F1-score over the CIC-BoT-IoT dataset. Our 3D-IDS outperforms AnomRank, the previous state-of-the-art method for anomaly detection on F1, by $15.27$ points over the EdgeIIoT dataset. We attribute the above results to the gains of our statistical disentanglement, representational disentanglement and dynamic graph diffusion method.
5.2.2. Comparisons of multi-classification.
@@
We compare the performance of our method to four baselines in declaring the specific attack type. These baselines include E-GraphSAGE (Lo et al., 2022), TGN (Rossi et al., 2020), ML (Wang et al., 2016), and AdaBoost (Schapire, 2013), as they are representative of different types of intrusion detection models and have been widely used in previous studies. Figure 4 shows that the proposed 3D-IDS consistently achieves the best results among all others over five datasets. For example, ours yields higher classification accuracy of up to $25$ points compared to the baseline TGN for detecting Injection attacks. The existing graph-based methods, including E-GraphSAGE, ML, and AdaBoost, perform inconsistently in the identification of complex attacks (e.g., MITM, Uploading, and XSS). We also observe that some attacks that are not easily detected by the baseline approaches, can be identified by the proposed 3D-IDS with high F1 scores. For example, E-GraphSAGE only achieves $18.34\%$ and $30.7\%$ F1 scores on CIC-ToN-IoT for MITM and Backdoor attacks, respectively, while our 3D-IDS is able to obtain higher than $23\%$ F1 score for each attack. These results further show the superiority of the proposed 3D-IDS. We also find that the average scores of our method on CIC-BoT-loT and CIC-ToN-loT are lower than the ones on the other three datasets. The underlying reason is the unbalanced attack distributions in the training set, where the dominant type may mislead the classifications. Such a finding aligns with previous work in the field of computer vision (Fisher et al., 2017). Nevertheless, the proposed 3D-IDS is still the best under such distributions. We leave this interesting observation as our future work.
5.2.3. Comparisons of unknown attacks.
@@
To further investigate the performance of detecting unknown attacks, we conduct experiments on the four attack types by discarding the corresponding instances in the train set and detecting them in the test set. We run our experiments $5$ times and report the mean and variance values with different random seeds. Table 2 reports the classification results on the CIC-ToN-IOT dataset. It shows that the statistical rule-based method Logistic Regression can only achieve as low as a $1.68$ F1 score for DDoS attacks, this confirms our analysis at the very beginning that rule-based methods can hardly detect unknown attacks. The score of graph-based E-GraphSAGE is much smaller than 3D-IDS, e.g., $6.05\%$ for MITM, indicating the limitations of the static graph in detecting unknown attacks. We also observe that TGN performs better than E-GraphSAGE, although both of them are graph-based methods. We attribute the improvement to the dynamic module for TCN. Nevertheless, our 3D-IDS outperforms all these methods by a large margin, with an average score of $33.65\%$ on the four attacks. The results also suggest that our method is more consistent in detecting various unknown attacks, showing the effectiveness of the two disentanglements.
Table 2. Unknown classification on the CIC-ToN-IoT dataset with metric F1-score (%)
Method Logistic Regression MSteam E-GraphSAGE TGN Ours (3D-IDS)
@@
5.3. Ablation Study
In this section, we conduct an ablation study on the CIC-ToN-IoT dataset to evaluate the effectiveness of each component. We remove our statistical disentanglement and denote it as ”w/o SD”. We use ”w/o RD” and ”w/o MLGRAND” to refer to the model that removes representational disentanglement and the multi-Layer graph diffusion module, respectively. Table 3 reports the comparison results. It shows that removing the multi-Layer graph diffusion module leads to the most significant performance degradation, e.g., an $18.33$ points decrease in AUC, indicating that it is the key component for the accuracy of the proposed 3D-IDS. Our second disentangled memory is also non-trivial to the overall detection accuracy, as removing this component can decrease the performance by $12.47$ points in AUC. We observe that the multi-layer graph diffusion module also benefits the model performance. The above ablation study further confirms the effectiveness of the three key components.
Table 3. Ablation study of 3D-IDS. F1-score (%), Precision (%), ROC-AUC (%), Recall (%), all metrics above are the average of five repeated experiments on the CIC-ToN-IoT dataset.
@@ 5.3. Ablation Study
3D-IDS(ours) 97.78$\pm$0.32 98.06$\pm$0.43 97.92$\pm$0.26 96.04$\pm$0.25
5.4. Discussion
@@
Figure 5. Statistical disentanglement of traffic features.
@@
RQ2222: How does the representational disentanglement benefit traffic classification for a specific attack? To answer this question, we track several Injection attack data in the CIC-ToN-IoT dataset and obtain the representation of these data in 3D-IDS and E-GraphSAGE. Meanwhile, we calculate the average of the embedding of the benign data in the two models as the scaling ratio. As shown in Figure 6, the representation values of E-GraphSAGE are much closer to the normal, which increases the probability of identifying the Injection attack as normal. It illustrates that as nodes aggregate, the discrepancies in features become blurred. The representation values of anomalous nodes gradually converge to normal, leading to incorrect link predictions. Values of our model in each dimension still deviate significantly from the normal values since ours imposes restrictions on the aggregation of information. The result proves the effectiveness of our model in maintaining a disentangled structure during the aggregation process, ensuring the presence of discrepancies, and leading to more accurate traffic classification for a specific attack.
(a) 3D-IDS
(b) 3D-IDS
@@
Figure 6. The comparison of node representation of the Injection attack after graph aggregation of our 3D-IDS and E-GraphSAGE. The grey line presents benign data.
Figure 7. Linear relationship between the disentangled edge vector and the disentangled node representation.
@@
RQ3333: How does the multi-layer diffusion module perform effectively for intrusion detection? We have illustrated the principle of multi-layer diffusion in Section 4.3. In this part, we take the MITM attack as an example to illustrate the effectiveness of spatial-temporal in intrusion detection.
@@
Figure 8 (a) shows a deep learning-based NIDS. When a MITM attack occurs, it is difficult to detect the intrusion since the spatial and temporal information of those packets is not considered. There are also some methods that only consider a single aspect of spatial and temporal information, such as E-GraphSAGE and MStream. In this case, for example, E-GraphSAGE mainly focuses on the spatial relationship of the set nodes and extracts features from them. However, we observe that different streams have their own timestamps from Table 4, so the lack of temporal information makes it impossible to analyze the dynamic structural changes of the edge. Similarly, taking the temporal information as the only effect factor will also get incomplete characteristics that do not contain spatial information (IP address). Moreover, some methods that take both the spatial and temporal information into account, such as Euler, take the snapshot method to capture the feature of the flow which does not achieve the synchronous update for spatial and temporal information. As shown in Figure 8 (b), intuitively, we can quickly detect that UE6 is an intrusion device of layer 1 when the flow changes from $SW2-SW3$ to $SW2-UE6$ and $UE6-SW3$ considering $SW2$, $SW3$ are layer $2$ devices. Also, we have noticed the changes in dynamic graph structure with a multi-Layer graph diffusion module to realize spatio-temporal coupling and synchronous updating. Overall, 3D-IDS performs best among these baselines in detecting various attacks.
Table 4. TimeStamp and IP
@@
Figure 8. Spatial-temporal coupling in intrusion detection.
@@
RQ4444: How does the two-step disentanglement facilitate the explainability of 3D-IDS? For this question, we rely on Figure 6 (a) as an example to recover the possible traffic features of a password attack. First, since the original features are retained after the double disentanglement, we can find some feature values that deviate significantly from the normal values in the node embedding. As shown in Figure 7, due to the disentangled edge representation, there exists a linear relationship between the disentangled edge vector and the disentangled node representation. Therefore, based on this relationship, we observe that the deviated features ”Fwd Pkt Std” and ”Fwd Pkt Len Max” present a strong correlation with the password attack.
6. CONCLUSION